Which of the following = "(n+1)! \/ (n-2)!"

  1. "n(n-1)(n-2)"
  2. "(n+1)n(n-1)!"
  3. "n(n-1)(n-2)!"
  4. "(n + 1)n(n - 1)"
  5. None of the above
1
Expert's answer
2021-06-30T16:50:12-0400
"\dfrac{(n+1)!}{(n-2)!}=\dfrac{(n+1)(n)(n-1)(n-2)!}{(n-2)!}="

"=(n+1)n(n-1)"

4. "(n+1)n(n-1)"